# A search using GEO600 for gravitational waves coincident with fast radio bursts from SGR 1935+2154

### v1

@@ Abstract
The magnetar SGR 1935+2154 is the only known Galactic source of fast radio bursts (FRBs). FRBs from SGR 1935+2154 were first detected by CHIME/FRB and STARE2 in 2020 April, after the conclusion of the LIGO, Virgo, and KAGRA Collaborations’ O3 observing run. Here we analyze four periods of gravitational wave (GW) data from the GEO600 detector coincident with four periods of FRB activity detected by CHIME/FRB, as well as X-ray glitches and X-ray bursts detected by NICER and NuSTAR close to the time of one of the FRBs. We do not detect any significant GW emission from any of the events. Instead, using a short-duration GW search (for bursts $\leq$ 1 s) we derive 50% (90%) upper limits of $10^{48}$ ($10^{49}$) erg for GWs at 300 Hz and $10^{49}$ ($10^{50}$) erg at 2 kHz, and constrain the GW-to-radio energy ratio to $\leq 10^{14}-10^{16}$. We also derive upper limits from a long-duration search for bursts with durations between 1 and 10 s. These represent the strictest upper limits on concurrent GW emission from FRBs.
@@ 1 Introduction
frb are a class of extremely energetic radio transients which are theorized to be associated with neutron stars (Platts et al., 2019; Petroff et al., 2019; Bailes, 2022; Zhang, 2023). To date, thousands of fast radio bursts have been detected. The majority of these have been discovered using the Canadian Hydrogen Intensity Mapping Experiment (CHIME) telescope (Amiri et al., 2022) by the CHIME/FRB Collaboration (CHIME/FRB) (CHIME/FRB Collaboration et al., 2018) [^1] https://www.chime-frb.ca/voevents. Though the origins of FRBs remain unknown, their dispersion measure (DM) as observed by radio telescopes localizes them to extragalactic (and even cosmological) distances (Thornton et al., 2013; Cordes & Chatterjee, 2019).
@@
SGR 1935+2154, as the first (and at the time of writing, only) FRB source to be confidently associated with a specific neutron star progenitor, presents a unique opportunity to search for GWs when the source is localized to a particular compact object. Additionally, at $\sim 6.6$ kpc (Zhou et al., 2020), it is more than two orders of magnitude nearer to Earth than the next closest FRB, which has been localized to the nearby galaxy M81, 3.6 Mpc away (Bhardwaj et al., 2021; Kirsten et al., 2022).
The four periods of FRB activity from SGR 1935+2154 fell in between the O3 and O4 observing runs of the LVK, when the LIGO and Virgo detectors were offline [^3] https://observing.docs.ligo.org/plan/. Fortunately, GEO600 (Grote et al., 2004; Lueck et al., 2010; Affeldt et al., 2014; Dooley et al., 2016), a GW detector in Hannover, Germany that is operated by members of the LIGO Scientific Collaboration, was observing in Astrowatch mode (Grote & the LIGO Scientific Collaboration, 2010) and collecting GW data during all four periods. The CHIME/FRB events for three of the four periods occurred when GEO600 was in observing mode, while the fourth FRB occurred within minutes of when GEO600 was observing (see Sec. 3).
In this paper, we analyze GEO600 data to search for GW emission coincident with the four FRBs observed by CHIME/FRB from SGR 1935+2154. We conduct two searches for unmodeled GW transients: one targeted at short-duration bursts with $\mathcal{O}{\rm(second)}$ durations, and another aimed at long-duration bursts lasting from 1 to 10 seconds. Due to SGR 1935+2154’s proximity, the results constitute the most sensitive searches for GWs from FRB sources to date, despite GEO600’s lower sensitivity compared to LIGO and Virgo (see Fig. 1). We also search for GWs coincident with the two X-ray glitches and the X-ray burst peak observed by NICER and NuSTAR in the hours around the FRB on 2022 October 14 (Hu et al., 2024). The paper is organized as follows: in Sec. 2, we describe the electromagnetic (EM) observations of FRBs from SGR 1935+2154. Section 3 details our short- and long-duration searches for GWs, with results presented in Sec. 4. We discuss the implications of these findings and conclude in Sec. 5.
Figure 1: Amplitude spectral density of GEO600 on 2020 April 28 compared to those of LIGO Hanford, LIGO Livingston, and Virgo during O3 (Abbott et al., 2020). While at low frequencies GEO600’s sensitivity is substantially diminished compared to that of the larger detectors, the gap narrows at frequencies around 2 kHz, near the expected neutron star f-mode frequency.
@@ 2 Fast radio bursts from SGR 1935+2154
Since 2020 April, SGR 1935+2154 has had multiple periods of high activity leading to the emission of FRBs. On 2020 October 08, CHIME/FRB observed three FRBs from SGR 1935+2154 arriving within a few seconds (Good & CHIME/FRB Collaboration, 2020; Pleunis & CHIME/FRB Collaboration, 2020; Giri et al., 2023). CHIME/FRB and the Green Bank Telescope (GBT) observed FRBs again two years later on 2022 October 14, with a CHIME/FRB event surrounded by five GBT FRBs within 1.5 seconds (Dong & CHIME/FRB Collaboration, 2022; Maan et al., 2022; Giri et al., 2023). During the days around the FRBs on 2022 October 14, SGR 1935+2154 was undergoing a period of intense X-ray burst activity. This burst storm began on October 10 (Mereghetti et al., 2022; Palmer, 2022) and was monitored by various telescopes such as NICER, NuSTAR, and XMM-Newton (see, e.g., Hu et al. (2024); Ibrahim et al. (2024)), which detected hundreds of milliseconds- to seconds-long bursts of high energy photons. The X-ray burst rate peaked during a flare 2.5 hr ($\pm$1 min) before the FRB and then steadily decreased over the next hours (Hu et al., 2024). In addition, the high-cadence monitoring observations allowed accurate measurements of the spin rate of SGR 1935+2154 (nominally 0.308 Hz; Israel et al. (2016)). The evolution of the spin rate showed that SGR 1935+2154 underwent a spin-up glitch about 4.4 hr ($\pm$30 min) before the FRB and another spin-up glitch about 4.4 hr ($\pm$30 min) after the FRB, while the magnetar’s spin-down rate between these two glitches was about one hundred times higher than its normal rate (Hu et al., 2024). X-ray bursts were also detected by GECAM and HEBS (Wang et al., 2022) and Konus-Wind (Frederiks et al., 2022) arriving within the expected FRB dispersion time. In addition to the NICER and NuSTAR observations mentioned above (Enoto et al., 2022), Insight-HXMT (Li et al., 2022) also observed X-rays from SGR 1935+2154 during this active period, though at the time of the FRB, all three were occulted by the Earth. Finally, a fourth FRB was detected by CHIME/FRB on 2022 December 01 (Pearlman & CHIME/FRB Collaboration, 2022; Giri et al., 2023), accompanied by a faint hard X-ray signal detected by Fermi-GBM (Younes et al., 2022).
@@
Figure 2: Radio energy versus luminosity distance for the SGR 1935+2154 FRBs investigated in this work (dark orange, Giri et al. (2023)) and for 749 other public FRBs published by CHIME/FRB and others (Petroff et al., 2016; Rajwade et al., 2020; CHIME/FRB Collaboration et al., 2021) (blue). The FRB sample and the calculation of distances and radio energies is described in Principe et al. (2023) (with the exception of the FRBs studied in Abbott et al. (2023), for which we use the lower bound 90% distances from that analysis). Note that the radio energies from CHIME/FRB (derived from fluxes and fluences) should be interpreted as lower limits (CHIME/FRB Collaboration et al., 2021; Andersen et al., 2023). We show the radio energy of the most luminous SGR 1935+2154 FRB, FRB20200428D, scaled to larger distances.
These SGR 1935+2154 FRBs are not quite like the rest of the population of FRBs, as mentioned in Sec. 1. As shown in Nimmo et al. (2022), they exhibit characteristics very similar to the extragalactic FRBs, but are a few orders of magnitude less luminous. Whether the SGR 1935+2154 FRBs are in the tail of the same population or truly occupy a different part of the phase space remains an open question. For example, while the 2020 April FRB from SGR 1935+2154 was several orders of magnitude less energetic than most FRBs, it was three orders of magnitude brighter than the next brightest previously observed radio flare from a magnetar (CHIME/FRB Collaboration et al., 2020). Figure 2 shows the radio energy as a function of distance for the FRBs from SGR 1935+2154, alongside a sample of 749 FRBs from CHIME/FRB [^4] https://www.chime-frb.ca/repeaters (CHIME/FRB Collaboration et al., 2021), the FRBCAT (Petroff et al., 2016), and the 76 m Lovell telescope (Rajwade et al., 2020) as collected and described in Principe et al. (2023). This could suggest that the emission mechanism which produces FRBs from SGR 1935+2154 may be different from that which results in FRBs from cosmological distances. Despite this reduced brightness compared to the typical FRB population, SGR 1935+2154’s proximity as the only known Galactic FRB source means that it presents the most promising opportunity to date for multi-wavelength and multi-messenger studies of FRB emitters.
@@ 2.1 Models for coincident GW-FRB emission
Most models which predict GW emission from FRB progenitors assume a CBC association, such as during or after the final stages of the CBC inspiral (Wang et al., 2016; Yamasaki et al., 2018), long before the CBC merger through interactions of magnetospheres (Zhang, 2020), or other interactions of compact binaries with their environments (see Platts et al. (2019) for a review of FRB theory). Prior studies such as Abbott et al. (2023) have searched for GWs from these sources using targeted matched-filter analyses, aimed at CBC sources. Since SGR 1935+2154 is not in a compact binary (Chrimes et al., 2022) and has exhibited multiple periods of FRB activity, we do not expect CBC-like GW emission from this source. Instead, as a magnetar, we can focus on only a few possible emission mechanisms for GWs coincident with FRBs.
Since at least some FRBs originate from magnetars, theories have drawn connections between them and magnetar giant flares (Tendulkar et al., 2016; Margalit et al., 2020; Cehula et al., 2024), which are are thought to be powered by magnetic activity near the surface of a neutron star (Thompson & Duncan, 1996; Gaensler et al., 2005). These giant flares are rare but so energetic that GW emission may be detectable (Ioka, 2001; Corsi & Owen, 2011). Quasi-periodic oscillations in the X-ray tails of giant flares may also create GWs (Levin & van Hoven, 2011; Quitzow-James et al., 2017). While no giant flares from SGR 1935+2154 were detected during its periods of FRB activity, the coincident X-ray activity suggests a potential link in the provenance of the high-energy EM emission.
@@
Moreover, neutron star glitches, such as those from SGR 1935+2154 in 2022 October investigated in this work, may emit GWs potentially observable by current GW detectors (Prix et al., 2011; Warszawski & Melatos, 2012; Melatos et al., 2015). Previous limits were derived for the Vela pulsar (located at 290 pc; Dodson et al. (2003)) glitch in 2006, providing limits on the emitted GW energy of the order of 1045 erg (Abadie et al., 2011).
@@ 3 Search for gravitational waves
Using data from GEO600 we search for generic GW transients from SGR 1935+2154 around the times of four FRBs detected by CHIME/FRB. GEO600 is a dual-recycled Michelson interferometer with folded arms and takes astrophysical observations in the 40 Hz - 6 kHz frequency band when operating in Astrowatch mode (Grote & the LIGO Scientific Collaboration, 2010). Over the past two decades, it has pioneered several key technologies for GW detectors (Affeldt et al., 2014; Lough et al., 2021). GEO600 has lower sensitivity compared to the Advanced LIGO & Advanced Virgo detectors ($\sim 10^{-22}/\sqrt{\mathrm{Hz}}$ at 1 kHz, see Fig. 1), but has strengths in uptime. It continued taking observations during the initial period of the COVID-19 pandemic and subsequent LIGO and Virgo upgrades throughout 2020-2022, during which CHIME/FRB observed these FRBs from SGR 1935+2154.
Given the unknown nature of the emission mechanism of the FRBs, we search for generic transient gravitational-wave signals present in the GEO600 data using two unmodeled burst searches: PySTAMP (Macquet et al., 2021), targeted at long-duration bursts with lengths from 1-10 s, and X-Pipeline (Sutton et al., 2010; Was et al., 2012), for short-duration bursts lasting less than 1 s. Previous searches for GWs coincident with FRBs, such as the ones presented in Abbott et al. (2023), also considered a possible CBC origin for the GW emission; the non-compact binary nature of SGR 1935+2154 precludes the use of CBC matched-filter searches. Given that SGR 1935+2154 is a magnetar, we follow the previous GW magnetar study presented in Abbott et al. (2024) and employ PySTAMP to perform a long-duration search, which has not previously been used for GW FRB analyses. Additionally, prior searches have typically been restricted to coincidences with FRBs where data from at least two GW detectors is available. For these FRBs from SGR 1935+2154, only GEO600 was observing, so we employ a single-detector search. This limits our ability to veto candidates based on coherence between detectors and the amount of background that can be estimated, reducing the search sensitivity, but given the extraordinary nature of these FRBs, we determined that a single-detector search in GEO600 data was warranted.
@@
For the 2020 April 28, 2020 October 08, and 2022 October 14 FRBs, we perform a search within an “on-source” time window starting at 1200 s before the infinite-frequency arrival time (i.e., the time accounting for the frequency-dependent delay introduced by the dispersion measure) of the FRB, $t_{0}$, and ending 120 s after. On 2020 October 08, three bursts were detected by CHIME/FRB within 3 s; we use the first time, corresponding to the FRB with the highest fluence on that day, as our $t_{0}$. The FRB on 2022 December 01 occurred during a time when GEO600 was not taking data, having exited observing mode approximately six minutes before the FRB, at 22:01:09 UTC. GEO600 returned to observing mode 23 minutes later, at 22:23:59 UTC. To be consistent with the on-source window for the other FRBs, we analyze the 800 s period of data beginning 1200 s before the FRB and ending shortly before GEO600 exited observing mode. Due to the large uncertainties in FRBGW models (as described above in Sec. 2.1), we use a wide extended on-source window of $[-1200,120]$ s. This allows us to probe a broad parameter space while keeping the detector behavior relatively stationary. We also employ a compact $[-4,4]$ s search window in the short-duration search to probe the time immediately surrounding the FRB with higher sensitivity. In addition, we search for emission around the time of three X-ray events detected by NuSTAR and NICER on 2022 October 14. These events consisted of a spin-up glitch, a peak in the X-ray burst emission, and another spin-up glitch. Since the uncertainty in their times is greater than a minute, we only perform an extended-window search, targeting a symmetric $[-1000,1000]$ s window, subject to data availability. Table 1 summarizes the times and windows for which we perform searches.
We restrict our search to frequencies from 300 Hz to 4096 Hz, with the lower cutoff set by the low frequency sensitivity of GEO600 and the upper cutoff aiming to capture neutron star crustal f-modes, which are predicted to fall at approximately 2 kHz (Andersson & Kokkotas, 1996).

### v2

@@ Abstract
The magnetar SGR 1935+2154 is the only known Galactic source of fast radio bursts (FRBs). FRBs from SGR 1935+2154 were first detected by the Canadian Hydrogen Intensity Mapping Experiment (CHIME)/FRB and the Survey for Transient Astronomical Radio Emission 2 in 2020 April, after the conclusion of the LIGO, Virgo, and KAGRA Collaborations’ O3 observing run. Here, we analyze four periods of gravitational wave (GW) data from the GEO600 detector coincident with four periods of FRB activity detected by CHIME/FRB, as well as X-ray glitches and X-ray bursts detected by NICER and NuSTAR close to the time of one of the FRBs. We do not detect any significant GW emission from any of the events. Instead, using a short-duration GW search (for bursts $\leq$ 1 s) we derive 50% (90%) upper limits of $10^{48}$ ($10^{49}$) erg for GWs at 300 Hz and $10^{49}$ ($10^{50}$) erg at 2 kHz, and constrain the GW-to-radio energy ratio to $\leq 10^{14}-10^{16}$. We also derive upper limits from a long-duration search for bursts with durations between 1 and 10 s. These represent the strictest upper limits on concurrent GW emission from FRBs.
@@ 1 Introduction
frb are a class of extremely energetic radio transients that are theorized to be associated with neutron stars (Thornton et al., 2013; Petroff et al., 2019; Platts et al., 2019; Bailes, 2022; Zhang, 2023). To date, thousands of fast radio bursts have been detected. The majority of these have been discovered using the Canadian Hydrogen Intensity Mapping Experiment (CHIME) telescope (Amiri et al., 2022) by the CHIME/FRB Collaboration (CHIME/FRB) (CHIME/FRB Collaboration et al., 2018) [^1] https://www.chime-frb.ca/voevents. Though the origins of FRBs remain unknown (Lorimer et al., 2024), their dispersion measure (DM) as observed by radio telescopes localizes them to extragalactic (and even cosmological) distances (Lorimer et al., 2007; Chatterjee et al., 2017; Cordes & Chatterjee, 2019).
@@
SGR 1935+2154, as the first (and at the time of writing, only) FRB source to be confidently associated with a specific neutron-star progenitor, presents a unique opportunity to search for GWs when the source is localized to a particular compact object. Additionally, at $\sim 6.6$ kpc (Zhou et al., 2020), it is more than two orders of magnitude nearer to Earth than the next closest FRB, which has been localized to the nearby galaxy M81, 3.6 Mpc away (Bhardwaj et al., 2021; Kirsten et al., 2022).
The four periods of FRB activity from SGR 1935+2154 fell between the O3 and O4 observing runs of the LVK, when the LIGO and Virgo detectors were offline [^3] https://observing.docs.ligo.org/plan/. Fortunately, GEO600 (Grote et al., 2004; Lueck et al., 2010; Affeldt et al., 2014; Dooley et al., 2016), a GW detector in Hannover, Germany that is operated by members of the LIGO Scientific Collaboration, was observing in Astrowatch mode (Grote & the LIGO Scientific Collaboration, 2010) and collecting GW data during all four periods. The CHIME/FRB events for three of the four periods occurred when GEO600 was in observing mode, while the fourth FRB occurred within minutes of when GEO600 was observing (see Sec. 3).
In this paper, we analyze GEO600 data to search for GW emission coincident with the four FRBs observed by CHIME/FRB from SGR 1935+2154. We conduct two searches for unmodeled GW transients: one targeted at short-duration bursts with $\mathcal{O}{\rm(second)}$ durations, and another aimed at long-duration bursts lasting from 1 to 10 seconds. Due to SGR 1935+2154’s proximity, the results constitute the most sensitive searches for GWs from FRB sources to date, despite GEO600’s lower sensitivity compared to LIGO and Virgo (see Fig. 1). We also search for GWs coincident with the two X-ray glitches and the X-ray burst peak observed by NICER and NuSTAR in the hours around the FRB on 2022 October 14 (Hu et al., 2024). This paper is organized as follows. In Sec. 2, we describe the electromagnetic (EM) observations of FRBs from SGR 1935+2154. Section 3 details our short- and long-duration searches for GWs, with results presented in Sec. 4. We discuss the implications of these findings and conclude in Sec. 5.
Figure 1: Amplitude spectral density of GEO600 on 2020 April 28 compared to those of LIGO Hanford, LIGO Livingston, and Virgo during O3 (Abbott et al., 2020). While at low frequencies GEO600’s sensitivity is substantially diminished compared to that of the larger detectors, the gap narrows at frequencies around 2 kHz, near the expected neutron-star f-mode frequency.
@@ 2 Fast radio bursts from SGR 1935+2154
Since 2020 April, SGR 1935+2154 has had multiple periods of high activity leading to the emission of FRBs. On 2020 October 08, CHIME/FRB observed three FRBs from SGR 1935+2154 arriving within a few seconds (Good & CHIME/FRB Collaboration, 2020; Pleunis & CHIME/FRB Collaboration, 2020; Giri et al., 2023). CHIME/FRB and the Green Bank Telescope (GBT) observed FRBs again two years later on 2022 October 14, with a CHIME/FRB event surrounded by five GBT FRBs within 1.5 seconds (Dong & CHIME/FRB Collaboration, 2022; Maan et al., 2022; Giri et al., 2023). During the days around the FRBs on 2022 October 14, SGR 1935+2154 was undergoing a period of intense X-ray burst activity. This burst storm began on October 10 (Mereghetti et al., 2022; Palmer, 2022) and was monitored by various telescopes, such as NICER, NuSTAR, and XMM-Newton (see, e.g., Hu et al. 2024; Ibrahim et al. 2024), which detected hundreds of milliseconds- to seconds-long bursts of high-energy photons. The X-ray burst rate peaked during a flare 2.5 hr ($\pm$1 min) before the FRB and then steadily decreased over the next hours (Hu et al., 2024). In addition, the high-cadence monitoring observations allowed accurate measurements of the spin rate of SGR 1935+2154 (nominally 0.308 Hz; Israel et al. 2016). The evolution of the spin rate showed that SGR 1935+2154 underwent a spin-up glitch about 4.4 hr ($\pm$30 min) before the FRB and another spin-up glitch about 4.4 hr ($\pm$30 min) after the FRB, while the magnetar’s spin-down rate between these two glitches was about one hundred times higher than its normal rate (Hu et al., 2024). X-ray bursts were also detected by GECAM and HEBS (Wang et al., 2022) and Konus-Wind (Frederiks et al., 2022) arriving within the expected FRB dispersion time. In addition to the NICER and NuSTAR observations mentioned above (Enoto et al., 2022), Insight-HXMT (Li et al., 2022) also observed X-rays from SGR 1935+2154 during this active period, though at the time of the FRB all three were occulted by the Earth. Finally, a fourth FRB was detected by CHIME/FRB on 2022 December 01 (Pearlman & CHIME/FRB Collaboration, 2022; Giri et al., 2023), accompanied by a faint hard X-ray signal detected by Fermi-GBM (Younes et al., 2022).
@@
Figure 2: Radio energy versus luminosity distance for the SGR 1935+2154 FRBs investigated in this work (dark orange, Giri et al. (2023)) and for 749 other public FRBs published by CHIME/FRB and others (Petroff et al., 2016; Rajwade et al., 2020; CHIME/FRB Collaboration et al., 2021) (blue). The FRB sample and the calculation of distances and radio energies is described in Principe et al. (2023) (with the exception of the FRBs studied in Abbott et al. (2023), for which we use the lower bound 90% distances from that analysis). Note that the radio energies from CHIME/FRB (derived from fluxes and fluences) should be interpreted as lower limits (CHIME/FRB Collaboration et al., 2021; Andersen et al., 2023). We show the radio energy required to produce a flare as bright as that the brighest FRB from SGR 1935+2154, FRB20200428D, as a function of distance.
These SGR 1935+2154 FRBs are not quite like the rest of the population of FRBs, as mentioned in Sec. 1. As shown in Nimmo et al. (2022), they exhibit characteristics very similar to the extragalactic FRBs, but are a few orders of magnitude less luminous. Whether the SGR 1935+2154 FRBs are in the tail of the same population or truly occupy a different part of the phase space remains an open question. For example, while the 2020 April FRB from SGR 1935+2154 was several orders of magnitude less energetic than most FRBs, it was three orders of magnitude brighter than the next brightest previously observed radio flare from a magnetar (CHIME/FRB Collaboration et al., 2020). Figure 2 shows the radio energy as a function of distance for the FRBs from SGR 1935+2154, alongside a sample of 749 FRBs from CHIME/FRB [^4] https://www.chime-frb.ca/repeaters (CHIME/FRB Collaboration et al., 2021), the FRBCAT (Petroff et al., 2016), and the 76 m Lovell telescope (Rajwade et al., 2020) as collected and described in Principe et al. (2023). This could suggest that the emission mechanism which produces FRBs from SGR 1935+2154 may be different from that which results in FRBs from cosmological distances. Despite this reduced brightness compared to the typical FRB population, SGR 1935+2154’s proximity as the only known Galactic FRB source means that it presents the most promising opportunity to date for multiwavelength and multimessenger studies of FRB emitters.
@@ 2.1 Models for coincident GW-FRB emission
Most models which predict GW emission from FRB progenitors assume a CBC association, such as during or after the final stages of the CBC inspiral (Wang et al., 2016; Yamasaki et al., 2018), long before the CBC merger through interactions of magnetospheres (Zhang, 2020), or other interactions of compact binaries with their environments (see Platts et al. (2019) for a review of FRB theory). Prior studies such as Abbott et al. (2023) have searched for GWs from these sources using targeted matched-filter analyses, aimed at CBC sources. Since SGR 1935+2154 is not in a compact binary (Chrimes et al., 2022) and has exhibited multiple periods of FRB activity, we do not expect CBC-like GW emission from this source. Instead, as a magnetar, we can focus on only a few possible emission mechanisms for GWs coincident with FRBs. In particular, because GWs are induced by time-varying quadrupole moments, we review models which predict EM magnetar activity associated with such moments.
Since at least some FRBs originate from magnetars, theories have drawn connections between them and magnetar giant flares (Tendulkar et al., 2016; Margalit et al., 2020; Cehula et al., 2024), which are are thought to be powered by magnetic activity near the surface of a neutron star (Thompson & Duncan, 1996; Gaensler et al., 2005). These giant flares are rare but so energetic that GW emission may be detectable due to hydromagnetic coupling of the magnetic dipole to the mass quadrupole (Ioka, 2001; Corsi & Owen, 2011). Quasi-periodic oscillations in the X-ray tails of giant flares may also create GWs through torsional or Alfvén modes which alter the star’s quadrupole moment (Levin & van Hoven, 2011; Glampedakis & Jones, 2014; Quitzow-James et al., 2017). While no giant flares from SGR 1935+2154 were detected during its periods of FRB activity, the coincident X-ray activity suggests a potential link in the provenance of the high-energy EM emission.
@@
Moreover, neutron-star glitches, such as those from SGR 1935+2154 in 2022 October investigated in this work, may emit GWs potentially observable by current GW detectors (Prix et al., 2011; Warszawski & Melatos, 2012; Melatos et al., 2015). Previous limits were derived for the Vela pulsar (located at 290 pc; Dodson et al. (2003)) glitch in 2006, providing limits on the emitted GW energy of the order of 1045 erg (Abadie et al., 2011).
@@ 3 Search for gravitational waves
Using data from GEO600, we search for generic GW transients from SGR 1935+2154 around the times of four FRBs detected by CHIME/FRB. GEO600 is a dual-recycled Michelson interferometer with folded arms and takes astrophysical observations in the 40 Hz - 6 kHz frequency band when operating in Astrowatch mode (Grote & the LIGO Scientific Collaboration, 2010). Over the past two decades, it has pioneered several key technologies for GW detectors (Affeldt et al., 2014; Lough et al., 2021). GEO600 has lower sensitivity compared to the Advanced LIGO and Advanced Virgo detectors ($\sim 10^{-22}/\sqrt{\mathrm{Hz}}$ at 1 kHz, see Fig. 1), but has strengths in uptime. It continued taking observations during the initial period of the COVID-19 pandemic and subsequent LIGO and Virgo upgrades throughout 2020-2022, during which CHIME/FRB observed these FRBs from SGR 1935+2154.
Given the unknown nature of the emission mechanism of the FRBs, we search for generic transient gravitational-wave signals present in the GEO600 data using two unmodeled burst searches: PySTAMP (Macquet et al., 2021), targeted at long-duration bursts with lengths from 1 to 10 s, and X-Pipeline (Sutton et al., 2010; Was et al., 2012), for short-duration bursts lasting less than 1 s. Previous searches for GWs coincident with FRBs, such as the ones presented in Abbott et al. (2023), also considered a possible CBC origin for the GW emission; the non-compact binary nature of SGR 1935+2154 precludes the use of CBC matched-filter searches. Given that SGR 1935+2154 is a magnetar, we follow the previous GW magnetar study presented in Abbott et al. (2024) and employ PySTAMP to perform a long-duration search, which has not previously been used for GW FRB analyses. Additionally, prior searches have typically been restricted to coincidences with FRBs where data from at least two GW detectors is available. For these FRBs from SGR 1935+2154, only GEO600 was observing, so we employ a single-detector search. This limits our ability to veto candidates based on coherence between detectors and the amount of background that can be estimated, reducing the search sensitivity, but given the extraordinary nature of these FRBs, we determined that a single-detector search in GEO600 data was warranted.
@@
For the 2020 April 28, 2020 October 08, and 2022 October 14 FRBs, we perform a search within an “on-source” time window starting at 1200 s before the infinite-frequency arrival time (i.e., the time accounting for the frequency-dependent delay introduced by the dispersion measure) of the FRB, $t_{0}$, and ending 120 s after. This asymmetric window is motivated by the expectation that any potential GWs are likely generated from the interior of the magnetar, preceding FRB emission from the magnetosphere or beyond. On 2020 October 08, three bursts were detected by CHIME/FRB within 3 s; we use the first time, corresponding to the FRB with the highest fluence on that day, as our $t_{0}$. The FRB on 2022 December 01 occurred during a time when GEO600 was not taking data, having exited observing mode approximately six minutes before the FRB, at 22:01:09 UTC. GEO600 returned to observing mode 23 minutes later, at 22:23:59 UTC. To be consistent with the on-source window for the other FRBs, we analyze the 800 s period of data beginning 1200 s before the FRB and ending shortly before GEO600 exited observing mode. Due to the large uncertainties in FRBGW models (as described above in Sec. 2.1), we use a wide extended on-source window of $[-1200,120]$ s. This allows us to probe a broad parameter space while keeping the detector behavior relatively stationary. We also employ a compact $[-4,4]$ s search window in the short-duration search to probe the time immediately surrounding the FRB with higher sensitivity. In addition, we search for emission around the time of three X-ray events detected by NuSTAR and NICER on 2022 October 14. These events consisted of a spin-up glitch, a peak in the X-ray burst emission, and another spin-up glitch. Since the uncertainty in their times is greater than a minute, we only perform an extended-window search, targeting a symmetric $[-1000,1000]$ s window, subject to data availability. Table 1 summarizes the times and windows for which we perform searches.
We restrict our search to frequencies from 300 Hz to 4096 Hz, with the lower cutoff set by the low frequency sensitivity of GEO600 and the upper cutoff aiming to capture neutron-star crustal f-modes, which are predicted to fall at approximately 2 kHz (Andersson & Kokkotas, 1996).